Solve using dual simplex method


Minimize z 2x₁ + 2x₂ + 4x3


2x+3x2 + 5x3 2 2


Subject to 3x1 + x₂ + 7x3 <3


x1 + 4x₂ + 6x3 ≤ 5


1
Expert's answer
2021-07-27T10:26:33-0400


In order to apply the dual simplex method, convert minZ to maxZ:

"maxZ=-2x_1-2x_2-4x_3"

subject to

"2x_1+3x_2+5x_3\le2"

"3x_1+x_2+7x_3\le3"

"x_1+4x_2+6x_3\le5"

"x_1,x_2,x_3\ge0"


The problem is converted to canonical form by adding slack, surplus and artificial variables:

  •  as the constraint-1 is of type '"\le" ' we should add slack variable S1
  •  as the constraint-2 is of type '"\le" ' we should add slack variable S2
  •  as the constraint-1 is of type '"\le" ' we should add slack variable S3


After introducing slack variables:

"maxZ=-2x_1-2x_2-4x_3+0S_1+0S_2+0S_3"

"2x_1+3x_2+5x_3+S_1=2"

"3x_1+x_2+7x_3+S_2=3"

"x_1+4x_2+6x_3+S_3=5"

"x_1,x_2,x_3,S_1,S_2,S_3\ge0"



Since all "Z_j-C_j\ge0" and all "X_{Bi}\ge0" thus the current solution is the optimal solution:

"x_1=x_2=x_2=0"